To prevent dehydration, intravenous (IV) solution is given to the hospital patients. Commonly used are solutions containing sodium chloride and glucose. A 150.0ml IV solution contains 6.60g glucose (C6H1206). What is the molarity of this solution?


1
Expert's answer
2021-03-04T06:33:20-0500

M=(m/Mr)/V= [mol/L];

M=(6.60/180)/0.15=0.24 mol/L


Answer:  molarity of this solution 0.24 M